OpenStudy (anonymous):

A boat can travel 114 miles downstream in the same time that it can travel 24 miles upstream. If the boat can travel 23 miles per hour in still water, find the speed of the current

OpenStudy (anonymous):

So, we know that distance = speed x time, which means that time = distance divided by the time. Since the time for the boat to travel both ways is the same, we can write the equation: \[114\div(23-c) = 24\div(23+c)\] where c is the speed of the current. 23-c is the speed of the boat against the current, and 23+c is the speed of the boat with the current. All we have to do is eliminate the denominators and solve for c. So the equation should look like this: \[114(23)+114c=23(24)+24c\]Solve for c and you should find the the current is 15mph